OpenStudy (anonymous):

[2.06] Solve the equation 2 over 5 (x − 4) = 2x for x. −8 negative 1 over 2 1 −1

OpenStudy (anonymous):

x-4 = 5x x = 5x + 4 1 = 5 + 4/x -4 = 4/x x= -1